# Exploring dynamical quantum phase transition from pure states to mixed states through extended Su-Schrieffer-Heeger models

### v2

@@ Abstract
We investigate dynamic quantum phase transitions (DQPTs) in both pure and mixed states within the framework of the extended SSH model, specifically analyzing the SSH-3 and SSH-4 models, which exhibit different symmetries. We find that the SSH-3 model, characterized by a chiral-like point symmetry rather than true chiral symmetry, supports robust localized edge states associated with its topological properties. Our results show that DQPTs for pure states occur following a quench that crosses the topological transition, even with an open energy band gap. For mixed states, DQPT behavior is consistent at low temperatures, but significant changes are observed at high temperatures, resulting in the emergence of multiple critical times. In contrast, the SSH-4 model, which possesses chiral symmetry, allows for the analysis of two distinct energy spectrum configurations. We conclude that the occurrence of DQPTs for pure states in the SSH-4 model necessitates a quench from an initial state without a band gap while crossing the critical point of the topological transition, whereas DQPTs are absent for mixed states at elevated temperatures.
@@ I Introduction
The paper is organized as follows: In Sec. II, we introduce the generalized SSH model and derive the time-dependent state of the system after a quantum quench. In Secs. III and IV, we investigate the behavior of DQPTs for both pure and mixed states in the SSH-3 model. In Sec. V, we study the property of the DQPT in the SSH-4 model. We conclude in Sec. VI
@@
We explore a generalized version of the SSH model, referred to as the SSH-m model. The Hamiltonian for this model is expressed as follows:
@@
where $t_{1},\cdots,t_{m-1}$ represent intracell couplings, and $t_{m}$ denotes the intercell coupling. In contrast to the SSH model, the generalized SSH-m model reveals richer variety of topological phase transitions Anastasiadis et al. (2022); Lee et al. (2022). In the Hamiltonian (1), we have imposed the periodic boundary condition. Consequently, the Bloch Hamiltonian is obtained by
@@
Firstly we consider a quantum quench starting from $t_{3}^{i}$ to $t_{3}^{f}$. The initial state is prepared at the pure state $|u_{k\mu}^{i}\rangle$, which satisfying
$H_{k}(t_{1}^{i})|u_{k\mu}^{f}\rangle=\varepsilon_{k\mu}^{i}|u_{k\mu}^{i}\rangle,\mu=1,\cdots,m.$ (3)
The time-evolved state following the quench is then given by
@@
To calculate Eq. (4), we can write the initial state as a linear superposition of eigenstates of post-quench Hamiltonian, which yields
@@
$\lambda(t)=-\lim_{N\rightarrow\infty}\frac{1}{N}\ln{[\mathcal{L}(t)]}=-\int_{-\pi}^{\pi}\frac{dk}{2\pi}\ln{r_{k}^{2}(t)}.$ (10)
Since $r_{k}(t)$ is well-defined, the singularities of the rate function $\lambda(t)$ occur only for $r_{k}(t)=0$, which is known as the Fisher zeros of the Loschmidt amplitude. At the Fisher zeros of $\mathcal{G}_{k}(t)$, the associate phase $\phi_{k}(t)$ is ill-defined, which induces the DTOP. According to Berry’s theory, the total phase $\phi_{k}(t)$ consists of a geometric phase $\phi_{k}^{G}(t)$ and a dynamical phase $\phi_{k}^{dyn}(t)$, in which the latter is contributed from the accumulation of the energy. In our model, the dynamical phase for the pure state can be obtained by
@@
As a result, the Pancharatnam geometric phase following the quench is given by
$\phi_{k}^{G}(t)=\mathrm{Arg}[\mathcal{G}_{k}(t)]-\phi_{k}^{dyn}(t).$ (12)
@@
$\mathbb{G}(t)=\mathrm{Tr}[\rho_{0}U(t)]=\prod_{k}gL_{k}(t)=\prod_{k}\mathrm{Tr}[\rho_{0k}U_{k}(t)].$ (13)
It is important to emphasize that the generalized Loschmidt amplitude proposed here is not limited by the parallel transport condition Sjöqvist et al. (2000); Tong et al. (2004); Lang et al. (2018); Tang et al. (2024). In our model, the initial density matrices for every $k$ take the form
@@
Similarly, we can also express the generalized Loschmidt amplitude in the polar coordinate, i.e. $\mathbb{G}_{k}(t)=R_{k}(t)e^{i[\Phi_{k}^{dyn}(t)+\Phi_{k}^{G}(t)]}$, where the dynamical phase for the mixed state is defined as
@@
Figure 1: The energy spectra as functions of $t_{3}$ in the SSH-3 models for (a) with $t_{1}=t_{2}=1$, and (b) with $t_{1}=1,t_{2}=1.2$. In (a), the gaps of the energy spectra vanish at $t_{3}=t_{1}=t_{2}=1$.
Figure 2: (a) The rate functions in the SSH-3 models with the gap-closing point [see Fig. 1 (a)] for the quenches from different initial states, where the quenches are from $t_{3}^{i}=0.2$ to $t_{3}^{f}=0.8$. (b) The rate functions for different initial states, in which the quenches are from $t_{3}^{i}=0.2$ to $t_{3}^{f}=1.1$.
@@
Figure 3: (a)-(c) The Pancharatnam geometric phases in the SSH-3 model with the gap-closing point [see Fig. 1 (a)] for different initial states ($|u_{k1}^{i}\rangle$, $|u_{k2}^{i}\rangle$, and $|u_{k3}^{i}\rangle$ ) are shown in momentum-time space, respectively. The dynamical phase vortices marked by the white circles exhibit the critical wave vectors and times. (d)-(f) The corresponding transition probabilities $|p_{k\nu}|^{2}$ for different initial states. The quench paths are from $t_{3}^{i}=0.2$ to $t_{3}^{f}=1.1$ with fixed $t_{1}=t_{2}=1$.
Since two-band systems (i.e. SSH-2 model in our paper) have been widely studied, we consider the main results of the SSH-3 and SSH-4 models in the following.
III DQPTs for pure states in SSH-3 model
In this section, we consider a generalization of the SSH model hosting three lattice sites in a unit cell. Notably, for the 1D SSH-3 model, the spectrum of the bulk Hamiltonian is not symmetric concerning zero Anastasiadis et al. (2022), implying that the system does not possess chiral symmetry. The winding number is also not well-defined for the SSH-3 model He and Chien (2020). However, the bulk Hamiltonian exhibits an additional interesting symmetry called point chirality, i.e. $\Gamma_{k}H_{k}\Gamma_{k}^{{\dagger}}=-H_{\pi-k}$ with $\Gamma_{k}=\mathrm{diag}(1,-1,1)$. In Fig. 1, we show the energy spectra of the bulk Hamiltonians for varying $t_{3}$ with fixed inner-cell hopping parameters $t_{1}$ and $t_{2}$. The Hamiltonian exhibits the gap closing only when $t_{1}=t_{2}=t_{3}$, which exactly corresponds to the case of the SSH model.
III.1 DQPTs for the case with closing gap
Firstly, we examine the dynamics of DQPT in the scenario where a gap closing point is present, as shown in the energy spectra in Fig. 1 (a). According to numerous studies, the collapse of a band gap is crucial in the theory of DQPTs. When the gap reopens, the system undergoes a DQPT. In Figs. 2(a) and (b), we present the typical rate function results for quenches that do not cross the gap closing point and for those that do, respectively. It is evident that a DQPT occurs only for quenches that cross the gap closing point. Additionally, we observe that point chiral symmetry causes the rate functions to initially be identical for both the lowest ($|u_{k1}^{i}\rangle$) and the highest ($|u_{k3}^{i}\rangle$) bands.
Furthermore, we would like more information by investigating the Pancharatnam geometric phase. As shown in Fig. 3 (a) to (c), we display the Pancharatnam geometric phases in the $(k,t)$ plane for the quenches from different initial bands respectively, where the quench is from $t_{3}^{i}=0.2$ to $t_{3}^{f}=1.2$. It is observed that the dynamical vortices as the signal of DQPTs show at the critical wave vectors $k_{c}$ and the critical times $t_{c}$. Interestingly, we notice that for quenches from the initial states $|u_{k1}^{i}\rangle$ and $|u_{k3}^{i}\rangle$, there is one critical wave vector for $k>0$, but two critical wave vectors ($k>0$) for the quench from $|u_{k2}^{i}\rangle$. These critical wave vectors satisfy the following relation:
@@ III.1 DQPTs for the case with closing gap
as we specifically marked in the Fig. 3.
In two-band systems, the critical wave vector $k_{c}$ is closely related to the equal transition probabilities, i.e. $|p_{k1}|^{2}=|p_{k2}|^{2}=1/2$. Our three-level model also has a similar relation involving two energy bands whose gap collapses at the critical case $t_{1}=t_{2}=t_{3}$. In Fig. 3 (d) to (f), we show the transition probabilities as functions of $k$ for quenches corresponding to that for Fig. 3 (a) to (c), respectively. It is observed that for the quench from $|u_{k1}^{i}\rangle$, the critical wave vector $k_{c1}$ can be obtained by solving
$|p_{k1}|^{2}=|p_{k2}|^{2}.$ (20)
Additionally, the wave vector $k_{c2}$ is obtain by
$|p_{k2}|^{2}=|p_{k3}|^{2}$ (21)
for quenches from $|u_{k2}^{i}\rangle$ or $|u_{k3}^{i}\rangle$.
III.2 Surprising DQPTs for the case with gap
Figure 4: (a) The rate functions in the SSH-3 models with the gap all the time [see Fig. 1 (b)] for the quenches from different initial states, where the quenches are from $t_{3}^{i}=0.2$ to $t_{3}^{f}=1.1$. (b) The rate functions for different initial states, in which the quenches are from $t_{3}^{i}=0.2$ to $t_{3}^{f}=1.3$.
Figure 5: The value of $\mathrm{max}(|p_{k2}|^{2})-\mathrm{min}(|p_{k1}|^{2})$ in the SSH-3 model with the gap all the time [see Fig. 1 (b)] for quenches from the initial state $|u_{k1}^{i}\rangle$. Here, we fix the parameter $t_{1}=1$ and the initial parameter $t_{3}^{f}=0.2$ to see the critical lines, i.e. $\mathrm{max}(|p_{k2}|^{2})-\mathrm{min}(|p_{k1}|^{2})=0$, for different $t_{2}$ and $t_{3}^{f}$.
Now, we consider the DQPT where the gap of energy bands is always present, as shown in Fig. 1 (b). In previous studies, it was generally believed that an energy gap would not lead to a DQPT in the system. However, in the SSH-3 model, the situation is different. By fixing $t_{1}=1$ and $t_{2}=1.2$, we show the rate function for the quench from $t_{3}^{i}=0.2$ to $t_{3}^{f}=1.1$ in Fig. 4 (a). Compared with that in Fig. 2 (b), the smooth rate function reveals that the presence of an energy gap may result in the disappearance of DQPTs. However, when $t_{3}^{f}=1.3$, the rate function shows a cusp-like singularity point [see Fig. 4 (b)]. This implies that there is a critical value $t_{3c}$, and when the quench crosses this critical value ($t_{3}^{f}>t_{3c}$), a DQPT will occur even though the gap is still there.
To determine the critical value, we can utilize condition (20) [or (21)]. According to the results in 3 (d) to (f), we know that the condition ensuring (20) is satisfied is given by
@@ III.2 Surprising DQPTs for the case with gap
Clearly, when the post-quench parameter is set at the critical point, we have $\mathrm{min}(|p_{k1}|^{2})=\mathrm{max}(|p_{k2}|^{2})$. In Fig. 5, we display the value of $\mathrm{max}(|p_{k2}|^{2})-\mathrm{min}(|p_{k1}|^{2})$ for quench from the lowest band $|u_{k1}^{i}\rangle$ varies as $t_{2}$ and $t_{3}^{f}$ with fixed $t_{1}=1$ and $t_{3}^{i}=0.2$, where the critical value $t_{3c}$ is denoted by $\mathrm{max}(|p_{k2}|^{2})-\mathrm{min}(|p_{k1}|^{2})=0$. It can be observed that the critical value $t_{3c}$ is closely related to $t_{2}$, although their relationship is not linear. To be precise, the critical value is always slightly less than $t_{2}$ except $t_{2}=1$. When $t_{2}=1$, the system returns to the scenario of the vanishing energy gap described in Sec. III.1. This result once again verifies the conclusion of the previous section, namely that quenching through the topological transition leads to a DQPT. In fact, the results of this section can also be understood as the same conclusion: when $t_{1}\neq t_{2}$ ($t_{2}>t_{1}$), the SSH-3 model has a hidden topological transition point at $t_{3}=t_{2}$. A change in the number of edge states under open boundary conditions characterizes this topological transition point.
IV DQPT for mixed states in SSH-3 model

### v3

@@ Abstract
We investigate dynamical quantum phase transitions (DQPTs) in both pure and mixed states within the extended SSH model framework, focusing on the SSH-3 and SSH-4 variants, which differ in symmetry properties. The SSH-3 model, characterized by a chiral-like point symmetry rather than true chiral symmetry, supports robust localized edge states tied to its topological nature. Our results show that for pure states, DQPTs occur after quenches crossing the topological transition, even when the energy band gap remains open. For mixed states, DQPT behavior aligns with pure states at low temperatures but undergoes significant changes at higher temperatures, including the emergence of multiple critical times. In contrast, the SSH-4 model, which possesses chiral symmetry, features four distinct energy spectrum configurations. We find that pure-state DQPTs arise only when the quench starts from a gapless initial state and crosses the critical topological point. At finite temperature, mixed-state DQPTs persist at low temperatures only if the corresponding pure-state quench induces DQPTs, but they disappear at elevated temperatures. These findings elucidate the interplay between symmetry, topology, and temperature in governing DQPTs within generalized SSH models.
@@ I Introduction
The paper is organized as follows: In Sec. II, we introduce the generalized SSH model and derive the time-dependent state of the system after a quantum quench. In Secs. III, we investigate the behaviors of DQPTs for both pure and mixed states in the SSH-3 model. In Sec. IV, we study the properties of the DQPTs for both pure and mixed states in the SSH-4 model. We conclude in Sec. V
@@
[…]
@@
where $t_{1},\cdots,t_{m-1}$ represent the intracell couplings, and $t_{m}$ denotes the intercell coupling. In contrast to the SSH model, the generalized SSH-m model reveals a richer variety of topological phase transitions Anastasiadis et al. (2022); Lee et al. (2022); Martinez Alvarez and Coutinho-Filho (2019). By imposing the periodic boundary condition, we obtain the bulk Hamiltonian as
@@
We now consider the quantum quench by suddenly changing the intercell coupling parameter $t_{m}$. The system is initialized in a pure state $|u_{k\mu}^{i}\rangle$, which satisfies
$H_{k}(t_{m}^{i})|u_{k\mu}^{f}\rangle=\varepsilon_{k\mu}^{i}|u_{k\mu}^{i}\rangle,\mu=1,\cdots,m.$ (3)
At $t=0$, the intercell coupling parameter is changed from $t_{m}^{i}$ to $t_{m}^{f}$. The time-evolved state following the quench is then given by
@@
To calculate Eq. (4), we can write the initial state as a linear superposition of eigenstates of the post-quench Hamiltonian, which yields
@@
$\lambda(t)=-\lim_{N\rightarrow\infty}\frac{1}{N}\ln{|G(t)|^{2}}=-\int_{-\pi}^{\pi}\frac{dk}{2\pi}\ln{r_{k}^{2}(t)}.$ (10)
Since $r_{k}(t)$ is well-defined, the singularities of the rate function $\lambda(t)$ occur only for $r_{k}(t)=0$, which is known as the Fisher zeros of the Loschmidt amplitude. At the Fisher zeros of $\mathcal{G}_{k}(t)$, the associated phase $\phi_{k}(t)$ is ill-defined, which induces the dynamical topological phase transition Budich and Heyl (2016). According to Berry’s theory, the total phase $\phi_{k}(t)$ consists of a geometric phase $\phi_{k}^{G}(t)$, known as the Pancharatnam geometric phase, and a dynamical phase $\phi_{k}^{dyn}(t)$, in which the latter is contributed from the accumulation of the energy. In our model, the dynamical phase for the pure state can be obtained by
@@
$\mathcal{G}(t)=\mathrm{Tr}[\rho_{0}U(t)]=\prod_{k}\mathrm{Tr}[\rho_{0k}U_{k}(t)],$ (12)
where $U(t)$ is the time-evolution operator. It is important to emphasize that the generalized Loschmidt amplitude proposed here is not limited by the parallel transport condition Sjöqvist et al. (2000); Tong et al. (2004); Lang et al. (2018); Tang et al. (2024). In our model, the initial density matrices for every $k$ take the form
@@
Similarly, we can also express the generalized Loschmidt amplitude in the polar coordinate, i.e. $\mathcal{G}_{k}(t)=R_{k}(t)e^{i[\Phi_{k}^{dyn}(t)+\Phi_{k}^{G}(t)]}$, where the dynamical phase for the mixed state is defined as
@@
As a result, the Pancharatnam geometric phase for both the pure and mixed states following the quench can be obtained by
$\phi_{k}^{G}(t)=\mathrm{Arg}[\mathcal{G}_{k}(t)]-\phi_{k}^{dyn}(t).$ (18)
III DQPTs in SSH-3 model
Figure 1: Energy spectra as a function of the intercell coupling $t_{3}$ in the SSH-3 model for (a) with $t_{1}=t_{2}=1$ (inversion-symmetric case), and for (b) with $t_{1}=1,t_{2}=1.2$ (inversion-symmetric broken case). The solid lines denote the bulk states, and the dotted lines represent the edge states. In (a), the energy gaps close at point $t_{3}=t_{1}=t_{2}=1$.
@@ III DQPTs in SSH-3 model
Figure 2: (a)-(c) The Pancharatnam geometric phases in the SSH-3 model with inversion symmetry from different initial states ($|u_{k1}^{i}\rangle$, $|u_{k2}^{i}\rangle$, and $|u_{k3}^{i}\rangle$ ) are shown in momentum-time space, respectively. The quench paths are from $t_{3}^{i}=0.2$ to $t_{3}^{f}=1.1$ with fixed $t_{1}=t_{2}=1$. The dynamical vortices marked by in white circles exhibit the critical wave vectors and times. (d)-(f) The corresponding transition probabilities $|p_{k\nu}|^{2}$ for different initial states.
[…]
In Fig. 1, we depict the energy spectra as a function of the intercell coupling $t_{3}$ for systems with and without inversion symmetry. In the inversion-symmetric case, the system exhibits a band-gap closing only when $t_{1}=t_{2}=t_{3}$. In the inversion-symmetric broken case, the system presents a pair of edge states for $t_{1}<t_{3}<t_{2}$, and two pairs of edge states for $t_{3}>t_{2}$. Notably, since the energy gaps remain open throughout the inversion-symmetry-broken regime, the shift in the number of edge states is not driven by the quantum phase transition. We will next explore how the properties of DQPTs are influenced by these different symmetry scenarios.
III.1 DQPTs for the inversion-symmetric case
Figure 3: The rate functions in the SSH-3 models with inversion symmetry for the quenches from different initial states, where the quenches in (a) are from $t_{3}^{i}=0.2$ to $t_{3}^{f}=0.8$, and in (b) are from $t_{3}^{i}=0.2$ to $t_{3}^{f}=1.1$.
Firstly, we investigate the properties of DQPTs in the SSH-3 model with inversion symmetry, of which the energy spectra show a band-gap closing at the critical point $t_{3c}=1$ [see Fig. 1 (a)]. In this context, the quench across the critical point of the quantum phase transition is generally regarded as a key condition for the occurrence of DQPT, although some counterexamples have been reported in certain systems Vajna and Dóra (2014); Cao et al. (2022). It is also worth noting that while the Fisher zero theory has traditionally served as a powerful and straightforward tool to identify DQPTs, its application becomes challenging in our model with multiple bands ($m\geq 3$) due to the complexity of solving the equation $\mathcal{G}_{k}(z)=0,z=\tau+it$. To circumvent this difficulty, we instead employ the concept of the Pancharatnam geometric phase, which reveals dynamical vortices at the critical momenta and critical times [see Figs.2 (a)-(c)].
Since the bulk Hamiltonian of the SSH-3 model possesses three pure eigenstates, $|u_{k\mu}\rangle$ with $\mu=1,2,3$, we analyze quenches starting from different initial pure states. Figures 3 display typical results of the rate functions corresponding to these quenches. In Figs. 3 (a), where the quenches do not cross the critical point $t_{3c}$, the rate functions remain smooth. In contrast, Fig. 3 (b) shows that the rate functions for quenches crossing $t_{3c}$ exhibit the cusp-like singularities at the critical time. This indicates that, in the inversion-symmetric system, the emergence of DQPT requires the quench to cross the critical point. Meanwhile, we observe that the rate functions for quenches starting from $|u_{k1}^{i}\rangle$ and $|u_{k3}^{i}\rangle$ coincide, which can be attributed to the underlying point chiral symmetry of the model.
Furthermore, we can obtain more information on the influence of the point chiral symmetry by analyzing details of the Pancharatnam geometric phases and the transition probabilities. Figs. 2 (a)-(c) depict the Pancharatnam geometric phases in the ($k,t$) plane for quenches from different initial states, and their corresponding transition probabilities $|p_{k\nu}|^{2}$ with $p_{k\nu}=\langle u_{k\nu}^{f}|u_{k\mu}^{i}\rangle$ are presented in (d)-(e). First of all, we find that the cases of quenches from $|u_{k1}^{i}\rangle$ and $|u_{k3}^{i}\rangle$ have one pair of critical wave vectors (±$\pm k_{c1}$ and ±$\pm k_{c2}$), and the case of quench from $|u_{k2}^{i}\rangle$ has two pairs of critical wave vectors. These critical wave vectors correspond to the same critical time and satisfy the following relation:
@@ III.1 DQPTs for the inversion-symmetric case
which matches the property of point chiral symmetry perfectly. Next, we notice that the critical wave vectors are exactly the intersection of the transition probabilities. For instance, in the case of quench from $|u_{k1}^{i}\rangle$, the lines $|p_{k1}|^{2}$ and $|p_{k2}|^{2}$ intersect at ±$\pm k_{c1}$; in the case of quench from $|u_{k3}^{i}\rangle$, the lines $|p_{k2}|^{2}$ and $|p_{k3}|^{2}$ intersect at ±$\pm k_{c2}$. This reminds us that, in two-band models, the critical wave vector $k_{c}$ is also closely related to the equal transition probabilities Cao et al. (2024b), i.e. $|p_{k1}|^{2}=|p_{k2}|^{2}=1/2$. The criterion can now be extended to multiband systems: When a DQPT occurs, the critical wave vector can be obtained by the intersection between two transition probabilities that correspond to the two neighbor energy bands.
III.2 DQPTs for the inversion-symmetric broken case
Figure 4: The rate functions in the SSH-3 models with broken inversion symmetry for the quenches from different initial states, where the quench in (a) is from $t_{3}^{i}=0.2$ to $t_{3}^{f}=1.1$ and in (b) is from $t_{3}^{i}=0.2$ to $t_{3}^{f}=1.3$.
Figure 5: The value of $\mathrm{max}(|p_{k2}|^{2})-\mathrm{min}(|p_{k1}|^{2})$ in the SSH-3 model with broken inversion symmetry for quenches from the initial state $|u_{k1}^{i}\rangle$. Here, we set the intracell coupling $t_{1}=1$ and the initial intercell coupling $t_{3}^{i}=0.2$. The critical line of the couplings $t_{2}$ and $t_{3}^{f}$ is denoted by the line $\mathrm{max}(|p_{k2}|^{2})-\mathrm{min}(|p_{k1}|^{2})=0$.
We now turn to the study of DQPTs in the SSH-3 model without inversion symmetry, where the energy bands remain gapped throughout [see Fig. 1 (b)]. Previous studies have generally held that the presence of an energy gap prevents the occurrence of DQPTs in such systems Huang and Balatsky (2016); Cao et al. (2024b). However, the SSH-3 model presents a distinct scenario. As the intercell coupling $t_{3}$ increases, the system experiences two discrete changes in the number of edge states. Notably, these changes do not correspond to phase transitions, since the energy gap remains finite at all times. This unique feature significantly influences the system’s nonequilibrium dynamical behavior.
In Fig. 4, we depict the rate functions for quenches starting from various initial states, with intracell couplings fixed at $t_{1}=1$ and $t_{2}=1.2$. In Fig. 4 (a), the quench path is from $t_{3}^{i}=0.2$ to $t_{3}^{f}=1.1$, crossing the first point at $t_{3}=1$ where the number of edge states changes. The resulting smooth rate functions indicate the absence of DQPTs in this regime. In contrast, Fig. 4 (b) shows a quench to $t_{3}^{f}=1.3$, which crosses the second such critical point at $t_{3}=1.2$. Here, the rate function displays a pronounced cusp-like singularity, signaling the occurrence of a DQPT. These findings suggest the existence of a critical value $t_{3c}$ such that when the quench crosses this threshold ($t_{3}^{f}>t_{3c}$), a DQPT can emerge even though the system remains gapped.
To determine the critical threshold $t_{3c}$ for the presence of DQPT, we can utilize the condition derived in the last section III.1: the critical wave vectors are exactly the intersection of the transition probabilities. For a quench starting from the initial pure state $|u_{k1}^{i}\rangle$, this condition can be expressed as
@@ III.2 DQPTs for the inversion-symmetric broken case
At the critical threshold $t_{3}^{f}$, this inequality becomes an equality:
$\mathrm{min}(|p_{k1}|^{2})=\mathrm{max}(|p_{k2}|^{2}).$ (21)
Furthermore, as illustrated in Fig. 4 (b), the critical threshold $t_{3c}$ is strongly related to the intracell coupling $t_{2}$. To explore this relationship, Fig. 5 shows the variation of $\mathrm{max}(|p_{k2}|^{2})-\mathrm{min}(|p_{k1}|^{2})=0$ as a function of $t_{2}$ and $t_{3}^{f}$ for quenches starting from $|u_{k1}^{i}\rangle$, with fixed parameters $t_{1}=1$ and $t_{3}^{i}=0.2$. The critical threshold $t_{3c}$ is identified by the contour line
$\mathrm{max}(|p_{k2}|^{2})-\mathrm{min}(|p_{k1}|^{2})=0.$ (22)
From the plot, it is evident that $t_{3c}$ increases with $t_{2}$, though the dependence is nonlinear. More specifically, $t_{3c}$ is generally slightly larger than $t_{2}$, except at the point $t_{2}=1$. At $t_{2}=1$, the system reduces to the SSH-3 with inversion symmetry, for which the threshold is exactly $t_{3c}=1$. This observation aligns well with previous results: in the inversion-symmetric regime, a DQPT occurs only when the quench crosses the system’s critical point.
III.3 DQPT for mixed states in SSH-3 model